How would you x so that x, x+2, and x+3 are consecutive terms of a geometric sequence and find the common ratio r of the sequence.
Answer by stanbon(75887)

x, x+2, and x+3 are consecutive terms of a geometric sequence and find the common ratio r of the sequence.
-------
Solve: (x+2)/x = (x+3)/(x+2)
------
x(x+3) = (x+2)^2
x^2+3x = x^2+4x+4
x = -4
----
Common ratio = -4
===================
cheers,
Stan H.